# Heavy Fermions as an Efficient Representation of Atomistic Strain and Relaxation in Twisted Bilayer Graphene

### v1

@@ I Introduction
Flat bands have figured prominently in the theoretical study of magic angle twisted bilayer graphene ever since their prediction [1, 2, 3, 4]. Most notably, the appearance of correlated insulators in their phase diagram can be explained from generalized models of flat band ferromagnetism, where exact Slater determinant groundstates and their excitations become computable in the perfectly flat, isolated limit [5, 6, 7, 8]. However, this strong coupling approach faces shortcomings in comparison to experiment [9, 10, 11, 12], where the ubiquitous effects of strain and lattice relaxation increase the dispersion of the flat bands and decrease their gap to remote bands [13, 14, 15, 16, 17, 18]. The low energy manifold of states appearing in this limit is split by strain and relaxation, which in turn dictate which phase is selected at integer filling [16, 11, 19]. A more robust framework is required to describe the low energy degrees of freedom in the presence of these unavoidable perturbations.
@@
Motivated a multitude of features in the strong coupling excitation spectra, Ref. [20] derived a mapping between the low-energy moiré bands of the Bistritzer-MacDonald (BM) Hamiltonian and a topological heavy fermion model [4], featuring trivial, impurity-like $f$-modes with no kinetic energy coupled to two sets of Dirac fermions ($c$-modes). This mapping preserves all symmetries of the original Hamiltonian (including an emergent particle-hole symmetry [21] and global $U(4)\times U(4)$ limits [22, 8]), and is a unitary change of basis within the low-energy Hilbert space, including the flat bands as well as the low-lying states in the nearby dispersive bands. In this way, one obtains a generalized periodic Anderson model [23, 24]. This model numerically reproduces the large number of low-lying symmetry-broken correlated insulator groundstates that emerge in Hartree-Fock at integer fillings [25, 26, 27, 28, 29, 30, 31, 32]. Moreover, one finds that the $f$-modes dominate the Hartree-Fock order parameter, yielding reliable analytical approximations for the groundstate and the excitation bands. An intuitive picture of the cascade physics emerges from these calculations, with heavy $f$-modes polarizing at integer fillings on top of the light $c$-mode Fermi sea [33]. This picture is borne out by effective field theory and dynamical mean-field theory calculations [34, 35, 36, 37, 38] and is compatible with experiment [39, 40, 41]. The analogy to Kondo physics [42, 43, 44] has opened a new direction in the study of TBG and beyond [45], especially away from integer filling.
@@
However, the interger-filling groundstates of the projected BM model are not robust to strain, which disrupts the approximate $U(4)\times U(4)$ symmetry that organizes the strong coupling groundstates in the flat band limit [5, 6, 7, 8] and allows competing states near the low-energy manifold, such as the Incommensurate Kekule Spiral (IKS)[16, 18, 46], to be selected. As we will show in this work, the expanded Hilbert space of the HF model easily accommodates strain and relaxation while preserving the heavy-fermion character of the bands. We make the case that heavy fermions provide a unifying formalism for TBG (and its multi-layer extensions [47, 48, 49]) which survives the inevitable perturbations in experiments [50, 51, 11, 52, 53]. The development of a single-particle Heavy Fermion (HF) model[20] including strain and relaxation is the focus of this paper. We study the interacting problem in coming work [54].
@@
Both strain and relaxation can be incorporated in the BM model. Early work [13, 15, 16, 18, 55, 56, 14, 16] approached these effects with a minimalist approach coupling strain as a pseudo-gauge field to the Dirac fermions [57] and implementing relaxation through renormalized parameters and $\mathbf{k}$-dependent (so-called “non-local”) interlayer tunneling. More recently, Kang and Vafek have shown that a generalized BM model can be derived systematically in a gradient expansion given a moiré relaxation profile [58], possibly including strain [59]. This formalism yields a moiré model that perfectly matches atomistic calculations, albeit requiring many parameters beyond the typical BM model — as expected in a microscopic Hamiltonian. Here we show that the HF model can quantitatively reproduce the Kang-Vafek (KV) model with only a handful of parameters (four in a minimal model), yielding a simple yet microscopically accurate Hamiltonian amenable to theoretical study.
@@ III Heterostrain and Non-local Tunneling in the BM Model
$\mathbf{\tilde{K}}^{l}$ $=\big{(}R(-l\theta/2)-\mathcal{E}_{l}^{T}\big{)}\mathbf{K}_{G}$ (13)
@@
$\mathbf{\tilde{q}}_{j}$ $=\big{(}R(\theta/2)-R(-\theta/2)\big{)}C^{j-1}_{3}\mathbf{K}_{G}-\mathcal{E}^{T}C^{j-1}_{3}\mathbf{K}_{G}$ (14)
where $\mathcal{E}=\mathcal{E}_{t}-\mathcal{E}_{b}$. Thus we see that heterostrain, where $\mathcal{E}_{t}=-\mathcal{E}_{b}$, couples directly to the moiré lattice, whereas homostrain $\mathcal{E}_{t}=+\mathcal{E}_{b}$ is absent. From here forward, we focus on heterostrain $\mathcal{E}$. Clearly $\mathcal{E}$ preserves $C_{2z}\mathcal{T}$ since it is uni-axial, but generically breaks $C_{3}$ and $C_{2x}$.
@@ IV Strain and Relaxation Corrections to the Heavy Fermion Model
$\delta h_{\epsilon}=\begin{pmatrix}c\,\boldsymbol{\epsilon}\cdot\boldsymbol{\sigma}^{*}&c^{\prime}\boldsymbol{\epsilon}\cdot\boldsymbol{\sigma}&i\gamma^{\prime}\epsilon_{+}\sigma_{3}\\ &M^{\prime}\epsilon_{+}\sigma_{2}&c^{\prime\prime}\boldsymbol{\epsilon}\cdot(\sigma_{0},i\sigma_{3})\\ h.c.&&M_{f}\boldsymbol{\epsilon}\cdot\boldsymbol{\sigma}^{*}\\ \end{pmatrix}$ (21)
in terms of $\boldsymbol{\epsilon}=(\epsilon_{xy},\epsilon_{-})$ and the 6 strain couplings which we compute to be
$(c,c^{\prime},c^{\prime\prime})=(7011,-5017,3144)\text{ meV}$ (22)
$(M_{f},\gamma^{\prime},M^{\prime})=(4252,3351,-4578)\text{ meV}\ .$
@@
with the $P$-breaking orbital potentials $\mu_{1},\mu_{2}$ (we always set $\mu_{f}=0$) and velocities $v_{1},v_{2},v_{3}$ computed to be
[…]
We find that $v_{3}$ is an order of magnitude smaller and can be neglected. In fact, we will show in Sec. VI.1 that the velocities $v_{1},v_{2}$ can also be neglected without losing accuracy in the flat bands.
@@
Figure 3: Comparison of symmetry-breaking $\text{error}_{g}(\mathbf{k})$. $(a)$ and $(b)$ are for the case of $C_{3z}$ in the BM and HF model, respectively. $(c)$ and $(d)$ are for the case of $C_{2x}$ in the BM and HF model, respectively.
@@ V.2 Relaxed HF Model
BM: 1.05 -24.8 -3.7 -4.3 1.6 14.4 4.5 0.2 -0.4 7011 -5017 3144 4252 3351 -4578
0.95 7.6 1.1 -3.5 -1. 22.2 9.5 0.2 -0.5 5156 -5334 6145 8525 2865 -3476
1.0 10.5 -3.5 -3.6 -1.1 23.6 10. 0.2 -0.5 5357 -5379 6195 8461 3052 -3946
1.05 14.5 -8.5 -3.8 -1.2 23.8 9.8 0.2 -0.5 5519 -5412 6232 8399 3220 -4342
1.1 18.5 -13.8 -3.9 -1.2 24.7 9.9 0.2 -0.5 5648 -5434 6255 8339 3371 -4682
1.15 21.8 -19.2 -4. -1.2 29.1 11.9 0.2 -0.5 5747 -5442 6260 8279 3506 -4975
1.2 27.2 -24.9 -4.1 -1.3 28.4 11.2 0.2 -0.5 5817 -5451 6270 8226 3626 -5233
Table 1: Heavy Fermion parameters for the fully relaxed KV model. All valued reported are in meV except for the velocities $v,v^{\prime},v_{1},v_{2}$ which are in eV Å. For comparison, the HF parameters in the BM model (including in $\boldsymbol{\Lambda}(\mathbf{r})$) are shown in the first line.
@@ VI Analytical Expressions for the Flat Bands
$\mathcal{U}^{\dagger}\delta h^{K}_{\epsilon}\mathcal{U}$ $=\frac{1}{v^{2}k^{2}+\gamma^{2}}\Big{(}2c^{\prime\prime}vk\gamma(\epsilon_{-}\sin\theta-\epsilon_{xy}\cos\theta)\sigma_{0}$ (44)
@@
where $\Gamma_{1}=\sigma_{3}\cos\theta+\sigma_{2}\sin\theta,\Gamma_{2}=\sigma_{2}\cos\theta-\sigma_{3}\sin\theta$ obey $\{\Gamma_{i},\Gamma_{j}\}=2\delta_{ij}\sigma_{0}$. A first observation is that only $c^{\prime\prime},M^{\prime},$ and $M_{f}$ appear at first order in perturbation theory. The most important parameter is $M_{f}$, the symmetry-breaking mass coupling the two $f$-modes. Away from $\Gamma$, all terms projected to the flat bands except $M_{f}$ go to zero and
@@
$\tan\theta_{n}$ $=\frac{|\boldsymbol{\epsilon}|\sqrt{M^{2}+M^{\prime 2}\epsilon_{+}^{2}}+M\epsilon_{xy}-M^{\prime}\epsilon_{+}\epsilon_{-}}{M\epsilon_{-}+M^{\prime}\epsilon_{+}\epsilon_{xy}}$
@@ VI.1 Minimal Models
Figure 6: Minimal HF model keeping the $\mu_{1},\mu_{2}$ PH-breaking parameters, and only the $c^{\prime\prime},M_{f}$ strain parameters. With only these parameters, good quantitative agreement is still achieved in the flat bands, and the key features of the dispersive bands are retained. $(a)$ depicts the BM bands with strain and PH-breaking corrections, and $(b)$ shows the KV bands. In both cases $\epsilon=0.1\%$ strain is used.
@@ VII Conclusion
The HF model was originally proposed to circumvent these difficulties [20]. It provides a faithful model whose simple form is amenable to analytical calculation, and whose locality and low dimensionality (after truncating the $c$-modes) make it suitable for intensive numerics. The inclusion of realistic strain and relaxation poses no difficulty for the formalism, underscoring the advantage of heavy fermions. In coming work, we leverage this result to develop an analytical understanding of the IKS as a heavy fermion state.
VIII Acknowledgements
The authors are grateful for important conversations and collaborations with Yves Kwan, Antoine Georges, Xi Dai, Andy Millis, Francisco Guinea, Roser Valentí, Giorgio Sangiovanni, Tim Wehling, Qimiao Si, Jian Kang, Dmitri K. Efetov, Keshav Singh, Liam Lau, Daniel Kaplan, Ryan Lee, Gautam Rai, Lorenzo Crippa, and Michael Scheer. This work was supported by Office of Basic Energy Sciences, Material Sciences and Engineering Division, U.S. Department of Energy (DOE) under Contracts No. DE-SC0016239 (B. A. B.), DE-FG02-99ER45790 (P. C.) and DE-SC0012704 (A. M. T.). J. H.-A. is supported by a Hertz Fellowship, with additional support from DOE Grant No. DE-SC0016239. J. Y. is supported by the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program and through the Gordon and Betty Moore Foundation’s EPiQS Initiative (Grant No. GBMF11070). D.C. acknowledges the hospitality of the Donostia International Physics Center, at which this work was carried out, and acknowledges support by the Simons Investigator Grant No. 404513, the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program, the Gordon and Betty Moore Foundation’s EPiQS Initiative (Grant No. GBMF11070), Office of Naval Research (ONR Grant No. N00014-20-1-2303), Global Collaborative Network Grant at Princeton University, BSF Israel US foundation No. 2018226. H.H. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101020833).
@@ A.1 Original BM Model
$\mathbf{q}_{1}$ $=R(\theta/2)\mathbf{K}_{G}-R(-\theta/2)\mathbf{K}_{G}=k_{\theta}(0,1),\quad\mathbf{q}_{j+1}=C_{3z}\mathbf{q}_{j},\quad k_{\theta}=\frac{4\pi}{3a}\times 2\sin\frac{\theta}{2}\ .$ (47)

### v3

@@ I Introduction
Flat bands have figured prominently in the theoretical study of magic angle twisted bilayer graphene ever since their prediction [1, 2, 3, 4]. Most notably, the appearance of correlated insulators in their phase diagram can be explained from generalized models of flat band ferromagnetism, where exact Slater determinant groundstates and their excitations become computable in the perfectly flat, isolated limit [5, 6, 7, 8]. However, this strong coupling approach faces shortcomings in comparison to experiment [9, 10, 11, 12], where the ubiquitous effects of strain and lattice relaxation increase the dispersion of the flat bands and decrease their gap to remote bands [13, 14, 15, 16, 17, 18, 19]. The low energy manifold of states appearing in this limit is split by strain and relaxation, which in turn dictate which phase is selected at integer filling [16, 11, 20]. A more robust framework is required to describe the low energy degrees of freedom in the presence of these unavoidable perturbations.
@@
Motivated a multitude of features in the strong coupling excitation spectra, Ref. [21] derived a mapping between the low-energy moiré bands of the Bistritzer-MacDonald (BM) Hamiltonian and a topological heavy fermion model [4], featuring trivial, impurity-like $f$-modes with no kinetic energy coupled to two sets of Dirac fermions ($c$-modes). This mapping preserves all symmetries of the original Hamiltonian (including an emergent particle-hole symmetry [22] and global $U(4)\times U(4)$ limits [23, 8]), and is a unitary change of basis within the low-energy Hilbert space, including the flat bands as well as the low-lying states in the nearby dispersive bands. In this way, one obtains a generalized periodic Anderson model [24, 25]. This model numerically reproduces the large number of low-lying symmetry-broken correlated insulator groundstates that emerge in Hartree-Fock at integer fillings [26, 27, 28, 29, 30, 31, 32, 33]. Moreover, one finds that the $f$-modes dominate the Hartree-Fock order parameter, yielding reliable analytical approximations for the groundstate and the excitation bands. An intuitive picture of the cascade physics emerges from these calculations, with heavy $f$-modes polarizing at integer fillings on top of the light $c$-mode Fermi sea [34]. This picture is borne out by effective field theory and dynamical mean-field theory calculations [35, 36, 37, 38, 39] and is compatible with experiment [40, 41, 42]. The analogy to Kondo physics [43, 44, 45] has opened a new direction in the study of TBG and beyond [46, 47], especially away from integer filling.
@@
However, the integer-filling groundstates of the projected BM model are not robust to strain, which disrupts the approximate $U(4)\times U(4)$ symmetry that organizes the strong coupling groundstates in the flat band limit [5, 6, 7, 8] and allows competing states near the low-energy manifold, such as the Incommensurate Kekule Spiral (IKS)[16, 18, 48], to be selected. As we will show in this work, the expanded Hilbert space of the HF model easily accommodates strain and relaxation while preserving the heavy-fermion character of the bands. We make the case that heavy fermions provide a unifying formalism for TBG (and its multi-layer extensions [49, 50, 51]) which survives the inevitable perturbations in experiments [52, 53, 11, 54, 55]. The development of a single-particle Heavy Fermion (HF) model[21] including strain and relaxation is the focus of this paper. We study the interacting problem in coming work [56].
@@
Both strain and relaxation can be incorporated in the BM model. Early work [13, 15, 57, 58, 16, 18, 59, 60, 14, 16] approached these effects with a minimalist approach coupling strain as a pseudo-gauge field to the Dirac fermions [61] and implementing relaxation through renormalized parameters and $\mathbf{k}$-dependent (so-called “non-local”) interlayer tunneling. More recently, Kang and Vafek have shown that a generalized BM model can be derived systematically in a gradient expansion given a moiré relaxation profile [62], possibly including strain [63]. This formalism yields a moiré model that perfectly matches atomistic calculations, albeit requiring many parameters beyond the typical BM model — as expected in a microscopic Hamiltonian. Here we show that the HF model can quantitatively reproduce the Kang-Vafek (KV) model with only a handful of parameters (four in a minimal model), yielding a simple yet microscopically accurate Hamiltonian amenable to theoretical study.
@@ III Heterostrain and Non-local Tunneling in the BM Model
$\mathbf{\tilde{K}}^{l}$ $=\big{(}R(l\theta/2)-\mathcal{E}_{l}^{T}\big{)}\mathbf{K}_{G}$ (13)
@@
$\mathbf{\tilde{q}}_{j}$ $=\mathbf{\tilde{K}}^{b}-\mathbf{\tilde{K}}^{t}$ (14)
which depends only on the heterostrain tensor
$\mathcal{E}=\mathcal{E}_{t}-\mathcal{E}_{b}$ (15)
which is parameterized by $\epsilon_{\pm},\epsilon_{xy}$ without the $l$ index. Thus we see that heterostrain, where $\mathcal{E}_{t}=-\mathcal{E}_{b}$, couples directly to the moiré lattice, whereas homostrain $\mathcal{E}_{t}=+\mathcal{E}_{b}$ is absent. From here forward, we focus on heterostrain $\mathcal{E}$. Clearly $\mathcal{E}$ preserves $C_{2z}\mathcal{T}$ since it is uni-axial, but generically breaks $C_{3}$ and $C_{2x}$.
@@ IV Strain and Relaxation Corrections to the Heavy Fermion Model
$\delta h_{\epsilon}=\begin{pmatrix}c\,\boldsymbol{\epsilon}\cdot\boldsymbol{\sigma}&c^{\prime}\boldsymbol{\epsilon}\cdot\boldsymbol{\sigma}^{*}&i\gamma^{\prime}\epsilon_{+}\sigma_{3}\\ &M^{\prime}\epsilon_{+}\sigma_{2}&c^{\prime\prime}\boldsymbol{\epsilon}\cdot(\sigma_{0},-i\sigma_{3})\\ h.c.&&M_{f}\boldsymbol{\epsilon}\cdot\boldsymbol{\sigma}\\ \end{pmatrix}$ (22)
in terms of $\boldsymbol{\epsilon}=(\epsilon_{xy},\epsilon_{-})$ and the 6 strain couplings which we compute in Table 1.
@@
with the $P$-breaking orbital potentials $\mu_{1},\mu_{2}$ (we always set $\mu_{f}=0$) and velocities $v_{1},v_{2},v_{3}$ (see Table 1). We find that $v_{3}$ is an order of magnitude smaller and can be neglected. In fact, we will show in Sec. VI.1 that the velocities $v_{1},v_{2}$ can also be neglected without losing accuracy in the flat bands.
@@
[…]
@@ V.2 Relaxed HF Model
BM: 1.05 -24.8 3.7 -4.3 1.6 14.4 4.5 0.2 -0.4 -8750 2050 -3362 4380 -3352 -4580
0.95 -7.6 1.1 -3.5 1. 22.2 9.5 0.2 -0.5 -5160 480 4660 -6510 -2410 -2560
1.0 -10.9 -3.5 -3.6 1.1 23.0 9.6 0.2 -0.5 -5750 550 4830 -6860 -2620 -3020
1.05 -14.5 -8.5 -3.8 1.2 23.8 9.8 0.2 -0.5 -6290 630 4990 -7190 -2810 -3430
1.1 -18.5 -13.8 -3.9 1.2 24.7 9.9 0.2 -0.5 -6760 710 5130 -7480 -2990 -3790
1.15 -22.9 -19.2 -4.0 1.3 25.5 10.1 0.2 -0.5 -7180 790 5270 -7740 -3150 -4110
1.2 -27.5 -24.9 -4.1 1.3 27.1 10.6 0.2 -0.5 -7540 880 5390 -7980 -3300 -4400
Table 1: Heavy Fermion parameters for the fully relaxed KV model. All valued reported are in meV except for the velocities $v,v^{\prime},v_{1},v_{2}$ which are in eV Å. For comparison, the HF parameters in the BM model (including $\boldsymbol{\Lambda}(\mathbf{r})$) are shown in the first line.
@@ VI Analytical Expressions for the Flat Bands
$\mathcal{U}^{\dagger}\delta h^{K}_{\epsilon}\mathcal{U}$ $=\frac{1}{v^{2}k^{2}+\gamma^{2}}\Big{(}-2c^{\prime\prime}vk\gamma(\epsilon_{-}\sin\theta+\epsilon_{xy}\cos\theta)\sigma_{0}$ (43)
@@
where $\Gamma_{1}=\sigma_{3}\cos\theta+\sigma_{2}\sin\theta,\Gamma_{2}=-(\sigma_{2}\cos\theta+\sigma_{3}\sin\theta)$ obey $\{\Gamma_{i},\Gamma_{j}\}=2\delta_{ij}\sigma_{0}$. A first observation is that only $c^{\prime\prime},M^{\prime},$ and $M_{f}$ appear at first order in perturbation theory. The most important parameter is $M_{f}$, the symmetry-breaking mass coupling the two $f$-modes. Away from $\Gamma$, all terms projected to the flat bands except $M_{f}$ go to zero and
@@
$\tan\theta_{n}$ $=\frac{|\boldsymbol{\epsilon}|\sqrt{M^{2}+M^{\prime 2}\epsilon_{+}^{2}}+M\epsilon_{xy}+M^{\prime}\epsilon_{+}\epsilon_{-}}{M^{\prime}\epsilon_{+}\epsilon_{xy}-M\epsilon_{-}}$
@@ VI.1 Minimal Models
Figure 6: Minimal HF model keeping the $\mu_{1},\mu_{2}$ PH-breaking parameters, and only the $c^{\prime\prime},M_{f}$ strain parameters. With only these parameters, good quantitative agreement is still achieved in the flat bands, and the key features of the dispersive bands are retained. $(a)$ depicts the BM bands with .$15\%.$ strain and PH-breaking corrections, and $(b)$ shows the KV bands with .$1\%.$ strain.
@@ VII Conclusion
The HF model was originally proposed to circumvent these difficulties [21]. It provides a faithful model whose simple form is amenable to analytical calculation, and whose locality and low dimensionality (after truncating the $c$-modes) make it suitable for intensive numerics. The inclusion of realistic strain and relaxation poses no difficulty for the formalism, underscoring the advantage of heavy fermions. In Ref. [56], we leverage this result to develop an analytical understanding of the IKS as a heavy fermion state.
VIII Acknowledgments
The authors are grateful for important conversations and collaborations with Yves Kwan, Antoine Georges, Xi Dai, Andy Millis, Francisco Guinea, Roser Valentí, Giorgio Sangiovanni, Tim Wehling, Qimiao Si, Jian Kang, Dmitri K. Efetov, Keshav Singh, Liam Lau, Daniel Kaplan, Ryan Lee, Gautam Rai, Lorenzo Crippa, and Michael Scheer. B. A. B. was supported by Office of Basic Energy Sciences, Material Sciences and Engineering Division, U.S. Department of Energy (DOE) under Contracts No. DE-SC0016239 and the Simons Collaboration on New Frontiers in Superconductivity (SFI-MPS-NFS-00006741-01). J. H.-A. is supported by a Hertz Fellowship, with additional support from DOE Grant No. DE-SC0016239. J. Y. is supported by the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program and through the Gordon and Betty Moore Foundation’s EPiQS Initiative (Grant No. GBMF11070). D.C. acknowledges the hospitality of the Donostia International Physics Center, at which this work was carried out, and acknowledges support by the Simons Investigator Grant No. 404513, the Gordon and Betty Moore Foundation through Grant No. GBMF8685 towards the Princeton theory program, the Gordon and Betty Moore Foundation’s EPiQS Initiative (Grant No. GBMF11070), Office of Naval Research (ONR Grant No. N00014-20-1-2303), Global Collaborative Network Grant at Princeton University, BSF Israel US foundation No. 2018226. H.H. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant Agreement No. 101020833).
@@ A.1 Original BM Model
$\mathbf{q}_{1}$ $=R(-\theta/2)\mathbf{K}_{G}-R(\theta/2)\mathbf{K}_{G}=-k_{\theta}(0,1),\quad\mathbf{q}_{j+1}=C_{3z}\mathbf{q}_{j},\quad k_{\theta}=\frac{4\pi}{3a}\times 2\sin\frac{\theta}{2}\ .$ (46)